71. A new pizza place, Fredrico’s opens in New York City. The average price of a medium pizza in New York is $10 and, because of large number of pizza sellers, this price will not be affected by the new entrant in the market. The owner of the Fredrico’s estimates that monthly total costs, including a normal profit will be

TC = 1000 + 2Q + 0.01Q2

To maximize total profit, how many pizzas should be produced each month? In the short run, how much economic profit the business will earn each month?
1
Expert's answer
2020-11-27T14:16:03-0500

The price of Pizza :

P = 10

Total Cost (TC) = 1000 + 2Q + 0.01Q2

Using the first order derivative of TC with respect to “Q” we can find the marginal cost:

"MC = \frac{dTC}{dQ} = 2 + 0.02Q"

Under perfect competitions the profit maximizing condition for Pizza hat is :

P = MC

10 = 2 + 0.02Q

8 = 0.02Q

"Q = \frac{8}{0.02} = 400"

Q = 400 pizzas (Profit maximizing quantity)

For calculating the economic profits we need the values of Total Revenue and Total cost:

Total Revenue:

"TR = P \times Q = 10 \times 400 = 4000"

Total Cost:

TC = 1000 + 2(400) + 0.01(400)2 = 1000 + 800 + 1600 = 3400

Economic Profits:

TR – TC = 4000 – 3400 = 600